OpenStudy (anonymous):

multiply and write the answer w/o exponents. (3x10^5) x (2x10^6)

OpenStudy (anonymous):

(3x10^5) x (2x10^6) = 3 x 2 x 10^5 x 10^6 = 6 x 10^11 = 600,000,000,000

OpenStudy (anonymous):

600,000,000,000